OpenStudy (anonymous):

Write the equation in standard form and slope-intercept form: Line Containg (4,-1) and having a slope of 5/2

OpenStudy (anonymous):

ok, so slope-intercept form is the form: y=mx+b where m is your slope and b is your y-intercept. so plug in your slope that you know already: y=(5/2)x + b now solve for b by plugging in the point you already have since (x,y) = (4,-1) we plug in: -1=(5/2)(4) + b and solve to get: -1=10+b -11=b so your equation is: y=(5/2)x-11

OpenStudy (anonymous):

Thanks for the help! Do you know how to write the equation in standard form?

OpenStudy (anonymous):

no prob! standard form is just the same thing but written with both x and y on the same side of the = so it would be y-(5/2)x=-11

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no prob, good luck with the homework